OpenStudy (anonymous):

The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line?

OpenStudy (anonymous):

y=(-(5/2)x) -5 y=((2/5)x)+2 y=((5/2)x )-2 y=-((2/5)x)-2